OpenStudy (brittany):

solve using quadratic formula with steps x^2-x-30=0

OpenStudy (anonymous):

I can solve it without the quadratic formula

OpenStudy (brittany):

oh how

OpenStudy (anonymous):

Well, there's three ways you can solve it without the quad formula: 1. Graph It 2. Complete The Square 3. Factor By Grouping

OpenStudy (brittany):

no thanks i need to learn quadratic formula

OpenStudy (anonymous):

Graphing is the easiest of course, but it requires a graphing calc

OpenStudy (anonymous):

Well, when using the quadratic formula, you always have to identify a, b, and c. Do you know how to do that?

OpenStudy (brittany):

yes

OpenStudy (brittany):

a=x^2 b=-x c= -30

OpenStudy (anonymous):

Good, that will help you

OpenStudy (anonymous):

Now, the only thing we have to do is know the quadratic formula and input the values of a, b, and c into it, then simplify

OpenStudy (brittany):

ok i usually have trouble simplfing it because idk wht step is first

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{b^2-4ac}}{2a} \]

OpenStudy (anonymous):

I see, well, we will plug in the values of a,b, and c first. You always use order of operations when simplifying expressions. The square root is the same as an exponent, so you have to evaluate that first.

OpenStudy (brittany):

in the square root should be x^2-4(x^2)(-30)

OpenStudy (anonymous):

Oh wait....your a, b, and c are incorrect

OpenStudy (anonymous):

a, b, and c, are numbers only that come before the variables in each term

OpenStudy (brittany):

oh?

OpenStudy (brittany):

x^2-x-30=0 a=1 b= -1 c = -30

OpenStudy (anonymous):

excellent

OpenStudy (anonymous):

Now what goes inside the square root?

OpenStudy (brittany):

-1^2-4(1)(-30)

OpenStudy (anonymous):

(-1)^2 -4(1)(-30) Good but make sure you put the -1 inside parentheses

OpenStudy (anonymous):

Because you will be squaring -1

OpenStudy (brittany):

which would make it 1

OpenStudy (anonymous):

Yes

OpenStudy (brittany):

1-4(1)(-30)

OpenStudy (anonymous):

excellent

OpenStudy (brittany):

now what

OpenStudy (anonymous):

Now you finish simplifying everything inside the square root

OpenStudy (brittany):

how?

OpenStudy (anonymous):

Order of operations still apply

OpenStudy (anonymous):

You multiply first before subtracting

OpenStudy (brittany):

multiply (1)(-30)?

OpenStudy (anonymous):

Multiply 4*1*-30

OpenStudy (brittany):

-120?

OpenStudy (anonymous):

yes, now you have 1-(-120)

OpenStudy (brittany):

121

OpenStudy (anonymous):

very good. Do you know how to take the square root of that?

OpenStudy (brittany):

well i normally wld use a calculater but i dont have one

OpenStudy (anonymous):

What is 11*11?

OpenStudy (brittany):

121

OpenStudy (anonymous):

Yes, so the square root of 121 is 11

OpenStudy (brittany):

ok so now its 1+-11/?

OpenStudy (anonymous):

\[\frac{1 \pm 11}{2*1}\]

OpenStudy (brittany):

1±11/2

OpenStudy (anonymous):

well, actually if you put it in that form, it isn't correct

OpenStudy (brittany):

oh?

OpenStudy (anonymous):

\[(1\pm 11)/2\]

OpenStudy (anonymous):

You have to make sure you put parentheses, that way we know that the whole numerator is divided by 2, not just 11/2

OpenStudy (brittany):

ok

OpenStudy (anonymous):

It's what happens when you use the linear form

OpenStudy (brittany):

how do i diivode tht by two

OpenStudy (anonymous):

If you write it as a top down fraction, you can leave it as: \[x = \frac{1 \pm 11}{2}\]

OpenStudy (anonymous):

Usually, you just leave it in that form because dividing by two would give decimals instead of whole numbers. We started with whole numbers so we should leave the answer in whole numbers

OpenStudy (brittany):

if i am aloud to use deciamls what do i do

OpenStudy (brittany):

i canuse deciamls rounded to the nearest tenth

OpenStudy (anonymous):

Then you can write it as: \[x = .5 \pm 5.5 \]

OpenStudy (brittany):

there positive?

OpenStudy (anonymous):

Or you can split it up into two values of x: x = -5 x = 6

OpenStudy (anonymous):

So in fact, we do end up with whole numbers, so you can put that as the answer

OpenStudy (brittany):

thank you so much i really appreciate it :)